# AI-Empowered Integrated Sensing and Communications

### v1

@@ IV-C Joint Beamforming and Waveform Design
Joint beamforming and waveform design can be approached with various objectives in mind. Maximizing the SNR is crucial for radar applications, while minimizing different types of interference (maximizing SINR) is essential for communications quality and transmission rate. In [96], the authors designed transmit precoders and receive combiners to minimize inter-user, intra-user, and radar-communications interference for ISAC systems. In [4], considering point and extended target scenarios, the authors proposed minimizing the CRB of radar sensing while ensuring a predefined level of SINR for each communication user.
@@ V-B Multi-Task Learning
While ISAC problems can be addressed using either single-task or multi-task learning, the latter is generally more suitable. As an alternative to single-task learning, we may train a single neural network for multiple task individually. Specifically, we may train the network to generate a waveform for $f_{\rm sens}(\cdot)$ and then retrain it for $f_{\rm comm}(\cdot)$. In this approach, a binary vector can be used as an input to distinguish the tasks, like in [114]. Even the radar component itself can involve multiple tasks. Traditional radar processing is modular, consisting of separate stages such as signal detection, target tracking, and classification. More specifically, radar processing typically involves sequential stages:
@@ V-C1 Plug-and-Play
An optimization algorithm may be too complex. In such a case, the plug-and-play approach can be applied to replace the most computation-intense components with a neural network [116], which is then trained on a set of similar optimizees, as previously described. This approach is especially useful when an optimization algorithm can be decomposed into update rules, allowing computationally intensive steps to be replaced with learnable neural networks.
@@ V-C2 Algorithm Unrolling
Algorithm unrolling is an alternative approach for model-based learning. It consists of three steps [117]. First, an iterative algorithm is picked/designed for solving an optimization problem. It is preferred that such iterative algorithm is simple, because in the second step, it is unrolled into a neural network. Unrolling stands for representation of computations in one iteration of an algorithm as a layer of a neural network, while the neural network is then build out of several such layers that correspond to the algorithm iterations. Third, a set of neural network parameters to be learned is selected. These may be, for example, step size parameters over the iterations of the corresponding optimization algorithm. An example is given in Section VI-B.
@@ V-D End-to-End Learning
In [134], an attention-based LSTM network was proposed for end-to-end predictive beamforming design in ISAC. This scheme captures the temporal correlation in reflected signal samples and directly determines the beamformer, offering a streamlined solution. To address high computational complexity caused by the need to continuously track vehicle motion parameters in vehicular ISAC, the authors in [135] designed an end-to-end predictive beamforming framework based on joint CRLBs. They evaluated sensing performance while ensuring an acceptable downlink communication sum-rate. Similarly, in [136], to adapt precoding policies to dynamic environments, the authors proposed an end-to-end learning method to directly learn downlink multi-user analog and digital hybrid precoders from the received uplink sounding reference signals. Additionally, they developed a parallel proactive optimization network to concurrently learn hybrid precoding.
@@ VI Representative Examples and Performance Evaluation
This section presents two learning-based ISAC waveform and beamformer designs. We describe the problem setups, explain the learning-based methods, and analyze their effectiveness through the following case studies:
@@
  • Case Study II utilizes a modified deep unrolling method to design hybrid analog/digital beamformers for ISAC.
@@ VI-A3 ISAC Signal Model
where $\mathbf{N}=[\mathbf{n}_{1},\cdots,\mathbf{n}_{\tau_{d}}]\in\mathbb{C}^{K\times\tau_{d}}$ is an Gaussian noise matrix having entries modeled as $\mathbf{n}_{i}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{M}),\forall i$. Next, we denote the desired received signal matrix for $K$ users by $\mathbf{Y}$. Then, we can rewrite (15) as [137]
@@ VI-A4 ISAC Waveform Optimization
    where $P_{T}$ denotes the ISAC waveform power per frame, and $\mathbf{X}_{0}$ is a reference radar waveform with desirable auto- and cross-correlation properties for target sensing [140, 137]. Specifically, $\mathbf{X}_{0}$ can be designed based on [137, Eq. (9)] for omnidirectional beampatterns and [137, Eq. (15)] for directional ones.
@@ VI-A5 A Novel Unsupervised Learning-based Technique
Output/Lambda Layer $2M\tau_{d}$
Here, we employ an unsupervised learning-based, single-task fully connected DNN to solve the optimization problem (22a)—a spacial form of (4)—and design the waveform. The DNN’s architecture is depicted in Fig. 5, and its structure is furnished in Table VII.
@@ VI-A6 Communication Performance Metrics
$\mathcal{R}_{k}=\sum^{K}_{k=1}\log[2]{1+\gamma_{k}},$ (25)
@@ VI-A7 Sensing Performance Metrics
VI-A8 Numerical Results and Performance Comparisons
@@
In our example, the carrier frequency is set to 3.2 GHz. The 3GPP urban microcell model [146, Table 5.1] is used to model the path loss, and the Rician factors are set to $K\in[1,3]$. The user-to-BS distance is set to be in $[50,200]$ m range. Unless otherwise specified, $P_{T}=1$ W, $M=\{16,32\}$, $K=4$, and $\tau_{d}=10$ are used. Moreover, a unit power QPSK modulation is adopted to generate the desired symbols in $\mathbf{D}$.
DNN implemented in Pytorch is used. We generate $10^{5}$ samples for the training, testing, and prediction stages. With $N=K(M+\tau_{d})+M\tau_{d}$, for $M=16$ the input layer has $5282N=$ 528 neurons, and the output layer has $M\tau_{d}=160$ neurons. About 60 % and 20 % of the total number samples are used during the offline training and testing phases, respectively. The remaining 20 % of generated samples are used to predict the learning-based ISAC waveforms during the online deployment phase. For training, Adam optimizer with the maximum epoch size of 500 and mini-batch size of 1000 is used. The initial learning rate for the optimizer is set to $0.001$, and the overfitting is circumvented by using early-stopping with patience 20.
@@
In Fig. 6, a performance comparison between optimization- and learning-based ISAC waveform designs in terms of the achievable sum rate is presented. The sum rate of the ISAC waveform designed by the proposed unsupervised learning-based approach is compared to that of four baseline techniques relying on optimization-based approaches [137]. The channel model from (11) is used, with Rician factors of 1.5, 2.7, 1.2, and 2.5 for the four users. The BS has $M=16$ antennas.
The genie-aided zero-MUI case (Case-I) is included in Fig. 6 to serve as a sum rate upper bound achieved using an ideal ISAC waveform. Case-D and Case-F shown in Fig. 6, we present the sum rate achievable by the proposed unsupervised learning-based technique for omnidirectional and directional desired beampatterns, respectively.
It can be unveiled from Fig. 6 that despite its sub-optimality, the learning-based ISAC waveform provides a comparable sum rate achievable by more complicated optimization-based counterparts. To be specific, by comparing Case-C with Case-E and Case-D with Case-F at a transmit SNR of 10 dB, we observe that the sum rate gaps between the learning-based and analytical baseline waveforms are as small as 0.876 % and 1.314 %, respectively. We also plot the sum rates achievable by the learning-based technique for both omnidirectional and directional ISAC waveforms (Case-G and Case-H) by increasing the number of BS antennas from 16 to 32 to demonstrate the sum rate gains and the scalability of the learning-based ISAC waveform design with respect to the number of BS antennas.
@@
Figure 7 shows the ROC curves of detection probability ($P_{D}$) versus false alarm probability ($P_{\rm FA}$) for target detection using both optimization- and learning-based ISAC waveform designs. The target detector proposed in [145, Eq. (69)] is used to compare the ROC performance. The ROC curves for Cases-2, 3 and 4 are plotted for the trade-off-based ISAC waveform designs obtained through the unsupervised learning-based approach. As a baseline comparison, in Case-1, the detector ROC curves with an ISAC waveform obtained via the SDR optimization-based solution for (22a)-(22b) in [137] are also plotted. For example, at $P_{\rm FA}=0.5$, the gap between Case-1 and Case-2 in terms of $P_{D}$ is about 0.0119.
Figure 7 also shows how $P_{D}$ can be improved for a given $P_{\rm FA}$ by prioritizing sensing over the achievable rate by suitably varying the weight factor ($\eta$) in the custom loss function (24). For example, by referring to Case-3 and Case-4 at $P_{\rm FA}=0.2$, we observe that $P_{D}$ can be improved by about 7.91 % by decreasing $\eta$ from 0.5 to 0.2. We also show in Fig. 7 that $P_{D}$ can be improved by using a larger array at the BS, compared in Case-2 with Case-3.
In Fig. 8, the transmit sensing beampatterns (19) of the learning-based ISAC waveform are compared with those of the optimization-based baseline techniques in [137]. We consider three target directions for the directional beampattern design as $-\pi/3$, 0, and $\pi/3$. The beampatterns provided by the unsupervised learning-based ISAC waveforms is closely align with the optimal beampatterns derived through SDR optimization-based technique for (22a)-(22b). This observation concludes that low-complexity learning-based approaches can be beneficial in circumventing the high computational complexity of optimization-based ISAC waveform designs.
VI-A9 Computational Complexity and Design Insights
The computational complexity of learning- and optimization-based ISAC waveforms can be compared in terms of floating-point operations. For the optimization-based ISAC waveform design with a sensing-centric omnidirectional beampattern, the complexity scales as $\mathcal{O}(MK\tau_{d}+M\tau_{d}^{2})$, due to two matrix multiplications and one SVD operation [137]. The complexity of the optimization-based ISAC waveform design with sensing-centric directional beampattern scales with $\mathcal{O}(M{\tau_{d}}^{2}+M^{2}\tau_{d}+MK\tau_{d}+M^{3}+M^{2}K)$ as it has one Cholesky decomposition, one SVD operation, and four matrix multiplications. The computational complexity of the trade-off based joint ISAC waveform design in (22a)-(22b) scales as $\mathcal{O}((M\tau_{d})^{3.5})$ when it is solved by leveraging SDR optimization-based technique in [142].
The computational complexity of a fully-connected layer scales as $\mathcal{O}((2D_{I}-1)D_{O})$, where $D_{I}$ and $D_{O}$ are the input and output dimensions, respectively [147]. In the worst-case scenario, they depend on the dimension of the input space $\mathcal{F}\in\mathbb{R}^{2N\times 1}$, where $N=K(M+\tau_{d})+M\tau_{d}$. During the ISAC waveform prediction stage, the overall computational complexity of the unsupervised learning-based technique scales as $\mathcal{O}(KM^{2}\tau_{d}+KM\tau^{2}_{d}+M^{2}\tau^{2}_{d})$.

### v2

@@ IV-C Joint Beamforming and Waveform Design
Joint beamforming and waveform design can be approached with various objectives in mind. Maximizing the SNR is crucial for radar applications, while minimizing different types of interference (maximizing SINR) is essential for communications quality and transmission rate. In [133], the authors designed transmit precoders and receive combiners to minimize inter-user, intra-user, and radar-communications interference for ISAC systems. In [4], considering point and extended target scenarios, the authors proposed minimizing the CRLB of radar sensing while ensuring a predefined level of SINR for each communication user.
@@ V-B Multi-Task Learning
While ISAC problems can be addressed using either single-task or multi-task learning, the latter is generally more suitable. As an alternative to single-task learning, we can train a single neural network with shared layers and some task-specific layers, as depicted in Fig. 2, to handle multiple tasks simultaneously. Despite it potential, multi-task learning-based ISAC has received considerably less attention in the literature compared to single-task learning ISAC. Representative examples can be found in [151], which jointly optimizes sensing, communication, and computation through a multi-objective formulation, solved using a multi-task learning model with a multiple-gradient descent algorithm. The goal is to simultaneously boost communication rate, enhance sensing precision, and reduce computing power consumption.
A multi-task learning architecture for channel estimation and position estimation in visible light integrated positioning and communication systems is proposed in [152]. The architecture consists of a shared network and two task-oriented sub-networks. Similarly, [153] investigates multi-task learning in communication and sensing, addressing tasks such as channel estimation, channel prediction, and human activity recognition. Their approach employs a Transformer-based architecture with self-supervised pretraining, followed by task-specific fine-tuning.
As an alternative to the multi-task learning structure in Fig. 2, which includes shared and task-specific layers, one may use only shared layers and distinguish tasks in a different way, e.g., by providing a binary vector, like in [115], as an additional input to indicate the task. In this approach, we can train the network to handle multiple tasks without explicit task-specific branches. Specifically, we may train a network to generate a waveform for $f_{\rm sens}(\cdot)$ and then retrain it for $f_{\rm comm}(\cdot)$. Even the radar component itself can involve multiple tasks. Traditional radar processing is modular, consisting of separate stages such as signal detection, target tracking, and classification. More specifically, radar processing typically involves sequential stages:
@@ V-C1 Plug-and-Play
When an optimization algorithm is too complex or computationally intensive, the plug-and-play (PnP) approach offers a practical alternative. In this framework, the algorithm is decomposed into modular update steps, and the most computationally demanding components, typically proximal operators or denoisers, are replaced with neural networks [33]. These networks are trained on a distribution of similar problems (optimizees), as previously discussed. This method is particularly effective when the structure of the original algorithm supports such modularity, allowing learning-based components to replace specific steps without changing the overall optimization logic.
To date, PnP methods have been primarily applied in the field of image reconstruction and have seen limited adoption in wireless communications or ISAC. Main reason may be that many PnP priors are designed for image processing and may not generalize well to communication tasks, where the statistical properties and physical constraints differ significantly. One notable exception is [155], where a deep PnP prior framework is used for multitask channel reconstruction in massive MIMO systems. In this approach, a shared regularization term acts as a learned wireless channel prior. However, the potential of PnP methods for ISAC applications remains largely unexplored and presents an open direction for future research. Unlike in imaging tasks, suitable PnP priors are still largely undeveloped. Therefore, while PnP may become useful in future ISAC scenarios, its applicability today remains limited due to the lack of domain-aware priors
@@ V-C2 Algorithm Unrolling
Algorithm unrolling is an alternative approach for model-based learning. It consists of three steps [34]. First, an iterative algorithm is picked/designed for solving an optimization problem. It is preferred that such iterative algorithm is simple, because in the second step, it is unrolled into a neural network. Unrolling stands for representation of computations in one iteration of an algorithm as a layer of a neural network, while the neural network is then build out of several such layers that correspond to the algorithm iterations. Third, a set of neural network parameters to be learned is selected. These may be, for example, step size parameters over the iterations of the corresponding optimization algorithm. There are many examples of algorithm unrolling applied to various communications and ISAC problems [19]. Such problems are typically non-convex, and this is the case when algorithm unrolling shows its power since unrolled (from an algorithm) and trained AI model is less affected by the existence of local minima, escapes them, and as a result finds a better suboptimal solution than an algorithm from which such AI model is unrolled. It also requires fewer layers in an unrolled network architecture than the number of iterations for an algorithm. However, the computations at each layer of an unrolled network are just equivalent to the computations at each iteration of an algorithm. Therefore, the computation saving for the pre-trained network come only from the reduced number of layers compared to the number of iteration of an algorithm, which is a factor reduction only. An illustrative example of algorithm unrolling for hybrid beamforming design problem, which have in detail discussions of the main points discussed above, can be found in Section VI-B.
@@ V-D End-to-End Learning
In [171], an attention-based LSTM network was proposed for end-to-end predictive beamforming design in ISAC. This scheme captures the temporal correlation in reflected signal samples and directly determines the beamformer, offering a streamlined solution. To address high computational complexity caused by the need to continuously track vehicle motion parameters in vehicular ISAC, the authors in [172] designed an end-to-end predictive beamforming framework based on joint CRLBs. They evaluated sensing performance while ensuring an acceptable downlink communication sum-rate. Similarly, in [173], to adapt precoding policies to dynamic environments, the authors proposed an end-to-end learning method to directly learn downlink multi-user analog and digital hybrid precoders from the received uplink sounding reference signals. In addition, they developed a parallel proactive optimization network to simultaneously learn hybrid precoding.
@@ VI Representative Examples and Performance Evaluation
This section presents three learning-based ISAC waveform, beamformer and constellation designs. We describe the problem setups, explain the learning-based methods, and analyze their effectiveness through the following case studies:
@@
  • Case Study II utilizes a modified deep unrolling method to design hybrid analog/digital beamformers.
  • Case Study III leverages autoencoders to perform end-to-end constellation design for ISAC.
@@ VI-A3 ISAC Signal Model
where $\mathbf{N}=[\mathbf{n}_{1},\cdots,\mathbf{n}_{\tau_{d}}]\in\mathbb{C}^{K\times\tau_{d}}$ is an Gaussian noise matrix having entries modeled as $\mathbf{n}_{i}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{K}),\forall i$. Next, we denote the desired received signal matrix for $K$ users by $\mathbf{D}$. Then, we can rewrite (13) as [62]
@@ VI-A4 ISAC Waveform Optimization
    where $P_{T}$ denotes the ISAC waveform power per frame, and $\mathbf{X}_{0}$ is a reference radar waveform with desirable auto- and cross-correlation properties for target sensing [176, 62]. Specifically, $\mathbf{X}_{0}$ can be designed based on [62, Eq. (9)] for omnidirectional beampatterns and [62, Eq. (15)] for directional ones.
@@ VI-A5 A Novel Unsupervised Learning-based Technique
Output/Lambda Layer $2M\tau_{d}$ -
Here, we employ an unsupervised learning-based, single-task fully connected DNN to solve the optimization problem (20a)—a special form of (2)—and design the waveform. The DNN’s architecture is depicted in Fig. 4, and its structure is furnished in Table VI.
@@ VI-A6 Communication Performance Metrics
$\mathcal{R}_{k}=\sum^{K}_{k=1}\mathrm{log}_{2}\left(1+\gamma_{k}\right),$ (23)
@@ VI-A7 Sensing Performance Metrics
VI-A8 Benchmarks for Performance Comparisons
The performance of the learning-based ISAC waveform designs is compared against the analytical benchmarks, which have been obtained through classical optimization theory. Our analytical benchmarks are elaborated below.
  1. I.
    Omni-directional sensing-centric waveforms: A typical sensing-centric waveform can be designed to minimize the MUI subject to a predefined beampattern (or equivalently transmit correlation matrix) as [62]
[…]
    subject to $\frac{1}{\tau_{d}}\mathbf{X}^{H}\mathbf{X}=\mathbf{C}_{\rm D},$ (25b)
    where $\mathbf{C}_{\rm D}$ is the desired transmit correlation matrix of the waveform. An omnidirectional beampattern is required for initial radar probing. For this case, we can set $\mathbf{C}_{\rm D}=P_{T}\mathbf{I}/M$ in (25b), where $P_{T}$ is the total transmit power. For an omnidirectional beampattern, the optimization problem in (25a)-(25b) can be solved optimally to obtain an analytical solution through SVD [62]. To this end, the ISAC transmit waveform $(\mathbf{X})$ satisfying (25a)-(25b) for an omnidirectional beampattern is computed in closed-form as reported in [62, Eq. (9)].
  2. II.
    Directional sensing-centric waveforms: A directional waveform is required when the radar is in its target tracking mode. When a directional beampattern is needed, a desired positive definite covariance matrix can be assigned to $\mathbf{C}_{\rm D}$ in (25b) by minimizing the mismatch between the desired beampattern and the obtained beampattern from (17) as reported in [177]. An analytical solution for an optional directional beampattern is obtained by solving the optimization problem in (25a)-(25b) by using SVD techniques as shown in [62, Eq. (15)].
  3. III.
    Sensing-communication trade-off based waveforms: Owing to the strict equality constraint (25b), the sensing-centric waveform design in (25a) may skew towards optimizing the sensing performance metrics. This design approach may hinder the achievable rate performance for the communication users. To mitigate such rate losses, in our numerical result comparisons, the fundamental sensing-communication trade-off is optimized to design the ISAC waveforms by solving the optimization problem in (20a)-(20b). To this end, the MUI for the communication users and the mismatch between the designed ($\mathbf{X}$) and reference ($\mathbf{X}_{0}$) ISAC waveforms have been minimized jointly to obtain an analytical solution for comparison purposes. We designed $\tilde{\mathbf{X}}$ for omnidirectional and directional beampatterns by following [62, Eq. (9)] and [62, Eq. (15)], respectively. Next, an analytical solution for the trade-off based ISAC waveform is obtained by first reformulating the optimization problem in (20a)-(20b) as a quadratically constrained quadratic program problem and then solving it through SDR techniques.
  4. IV.
    Ideal communication-centric waveforms: A genie-aided zero-MUI waveform is also used to obtain a communication performance upper bound. Specifically, the achievable rate at the $k$th user by using this zero-MUI is given by $\gamma_{k}=\mathbb{E}\left[|D_{kq}|^{2}\right]/\sigma^{2}$.
VI-A9 Numerical Results and Performance Comparisons
@@
In our example, the carrier frequency is set to 3.2 GHz. The 3GPP urban microcell model [182, Table 5.1] is used to model the path loss, and the Rician factors are set to $K\in[1,3]$. The user-to-BS distance is set to be in $[50,200]$ m range. Unless otherwise specified, $P_{T}=1$ W, $M=\{16,32\}$, $K=4$, and $\tau_{d}=10$ are used. Moreover, a unit power QPSK modulation is adopted to generate the desired symbols in $\mathbf{D}$.
DNN implemented in Pytorch is used. We generate $10^{5}$ samples for the training, testing, and prediction stages. With $N=K(M+\tau_{d})+M\tau_{d}$, for $M=16$ the input layer has $2N=528$ neurons, and the output layer has $M\tau_{d}=160$ neurons. About 60 % and 20 % of the total number samples are used during the offline training and testing phases, respectively. The remaining 20 % of generated samples are used to predict the learning-based ISAC waveforms during the online deployment phase. For training, Adam optimizer with the maximum epoch size of 500 and mini-batch size of 1000 is used. The initial learning rate for the optimizer is set to $0.001$, and the overfitting is circumvented by using early-stopping with patience 20.
@@
9.1) Comparison of Communication Performance: In Fig. 8, a performance comparison in terms of a communication metric is presented between the learning-based ISAC waveform designs and classical optimization based analytical waveform designs. Specifically, the achievable sum rate of the ISAC waveforms designed by the proposed unsupervised learning-based approach is compared to that of four analytical waveform benchmarks described in Section VI-A8. The channel model (9) is used, with Rician factors of 1.5, 2.7, 1.2, and 2.5 for the four users. The BS has $M=16$ or $M=32$ antennas. The genie-aided zero-MUI case (Case-I) is included in Fig. 8 to serve as a sum rate upper bound achieved using an ideal communication-centric ISAC waveform. Case-D and Case-F shown in Fig. 8, we present the sum rate achievable by the proposed unsupervised learning-based technique for omnidirectional and directional desired beampatterns, respectively.
It can be unveiled from Fig. 8 that despite its sub-optimality, the learning-based ISAC waveform provides a comparable sum rate achievable by more complicated optimization-based counterparts. To be specific, by comparing Case-C with Case-E and Case-D with Case-F at a transmit SNR of 10 dB, the sum rate gaps between the learning-based and analytical baseline waveforms are as small as 0.88 % and 1.31 %, respectively. Nevertheless, the classical optimization-based analytical waveform designs achieve this slight sum rate gain at the expense of significant computational complexity cost as elaborated in Section VI-A10.
We also plot the sum rates achievable by the learning-based technique for both omnidirectional and directional ISAC waveforms (Case-G and Case-H) by increasing the number of BS antennas from 16 to 32 to demonstrate the sum rate gains and the scalability of the learning-based ISAC waveform design with respect to the number of BS antennas.
@@
9.2) Comparison of Sensing Performance: In Fig. 6, a comparison of performance in terms of sensing metrics is given between the ISAC waveforms designed through the proposed learning-based and classical optimization-based techniques. Specifically, Fig. 6 shows the ROC curves of detection probability ($P_{D}$) versus false alarm probability ($P_{\rm FA}$) for target detection. As a benchmark comparison, in Case-1, the detector ROC curves with an ISAC waveform obtained by analytically solving the sensing-communication trade-off based design (20a)-(20b) elaborated in Section VI-A8 are also plotted. Moreover, the target detector proposed in [181, Eq. (69)] is used to compare the ROC performance. The ROC curves for Cases-2, 3 and 4 are plotted for the trade-off-based ISAC waveform designs obtained through the proposed unsupervised learning-based approach. Through our benchmark comparison, Fig. 6 unveils that the performance gap in terms of sensing metrics between the proposed learning-based and analytical-based ISAC waveform designs is again very small. For example, at $P_{\rm FA}=0.5$, the gap between Case-1 (analytical-based) and Case-2 (learning-based) in terms of $P_{D}$ is about 0.0119.
Figure 6 also shows how $P_{D}$ can be improved for a given $P_{\rm FA}$ by prioritizing sensing over the achievable rate by suitably varying the weight factor ($\eta$) in the custom loss function (22). For example, by referring to Case-3 and Case-4 at $P_{\rm FA}=0.2$, we observe that $P_{D}$ can be improved by about 7.91 % by decreasing $\eta$ from 0.5 to 0.2. We also show in Fig. 6 that $P_{D}$ can be improved by using a larger array at the BS, compared in Case-2 (with $M=16$ BS antennas) with Case-3 (with $M=32$ BS antennas). For example, at $P_{\rm FA}=0.2$, $P_{D}$ can be increased by about 18.88 % by doubling the number of BS antennas from 16 to 32.
@@
9.3) Comparison of Unified Beampatterns for ISAC: In Fig. 7, the unified ISAC transmit beampatterns (17) of the learning-based ISAC waveform are compared with those of the optimization-based analyical benchmarks presented in Section VI-A8. We consider three target directions for the directional beampattern design as $\pi/3$, 0, and $\pi/3$. Fig. 7 reveals that the unified ISAC beampatterns provided by the proposed unsupervised learning-based technique is closely aligned with the optimal analytical beampatterns derived through SDR optimization-based technique for (20a)-(20b). For both analytical and learning-based techniques, the maximum gains of the beampatterms appear in the chosen target directions ($\pi/3$, 0, and $\pi/3$). As further elaborated in Section VI-A10, the analytical-based ISAC waveform designs possess a considerable computational complexity over the learning-based alternatives. Hence, the observations in Fig. 7 conclude that low-complexity learning-based approaches can be beneficial in circumventing the high computational complexity of optimization-based ISAC waveform designs while achieving near-optimal unified ISAC beampatterns.
VI-A10 Computational Complexity Comparison
In this subsection, a computational complexity comparison is presented between the proposed learning-based and analytical optimization-based ISAC waveform design approaches. The underlying computational complexity can be compared in terms of floating-point operations. First, the computational complexity of analytical ISAC waveform designs, which have been presented as benchmarks in Section VI-A8, is presented as follows. For the optimization-based analytical ISAC waveform design with a sensing-centric omnidirectional beampattern, the complexity scales as $\mathcal{O}(MK\tau_{d}+M\tau_{d}^{2})$, due to two matrix multiplications and one SVD operation [62]. Since the underlying optimization problem can be cast as an orthogonal procrustes problem (OPP) [183], it converges to a globally optimal solution with a linear convergence rate. The complexity of the optimization-based ISAC waveform design with sensing-centric directional beampattern scales with $\mathcal{O}(M{\tau_{d}}^{2}+M^{2}\tau_{d}+MK\tau_{d}+M^{3}+M^{2}K)$ as it has one Cholesky decomposition, one SVD operation, and four matrix multiplications. Again, this optimization procedure relies on an OPP, and hence, it converges to a globally optimal solution with a linear convergence rate. The computational complexity of the trade-off based joint ISAC waveform design in (20a)-(20b) scales as $\mathcal{O}((M\tau_{d})^{3.5})$ when it is solved by leveraging SDR optimization-based technique in [178]. This SDR is tight, and its solution is rank-1, yielding a global optimizer. The corresponding low-complexity algorithm uses the Golden-section search, having a linear convergence rate [184].
Next, we evalute the computational complexity of the proposed learning-based ISAC waveform. The computational complexity of a fully-connected layer scales as $\mathcal{O}((2D_{I}-1)D_{O})$, where $D_{I}$ and $D_{O}$ are the input and output dimensions, respectively [185]. In the worst-case scenario, they depend on the dimension of the input space $\mathcal{F}\in\mathbb{R}^{2N\times 1}$, where $N=K(M+\tau_{d})+M\tau_{d}$. During the ISAC waveform prediction stage, the overall computational complexity of the unsupervised learning-based technique scales as $\mathcal{O}(KM^{2}\tau_{d}+KM\tau^{2}_{d}+M^{2}\tau^{2}_{d})$.
The above analysis reveals that the unsupervised learning-based ISAC waveform design approached proposed to solve (20a)-(20b) has a much lower computational complexity than that of the corresponding optimization-based solutions. Although suboptimal, the learning-based ISAC waveform designs can be useful as a computationally efficient alternative to more complicated optimization-based ISAC waveform designs, particularly when the system scales in terms of the numbers of BS antennas and the communication users.
VI-A11 Generalization Limitations
We analyze performance degradation in the proposed unsupervised learning-based unified ISAC waveform design due to generalization limitations. Particularly, two such limitations, channel changes/aging due to user mobility and system topology changes due to increments of the total number of users, are investigated as follows:
  1. I.
    Channel Aging: The ISAC waveform optimized for an aged channel $\mathbf{h}_{k}[n-1]$ is used to transmit data for a case with user mobility. In this case, the true channel $\mathbf{h}_{k}[n]$ differs from the aged channel $\mathbf{h}_{k}[n-1]$, and their relationship can be modeled as follows:
    $\!\!\!\!\!\!\!\!\!\mathbf{h}_{k}[n]$ $=$ $\sqrt{\frac{K_{h_{k}}\eta_{h_{k}}}{K_{h_{k}}+1}}\bar{\mathbf{h}}_{k}[n]+\sqrt{\frac{\zeta_{h_{k}}}{K_{h_{k}}+1}}\tilde{\mathbf{h}}_{k}[n],$ (26)
    where $\mathbf{h}_{k}[n]$ is a sampled version of (9). Under user mobility, the LoS component $\bar{\mathbf{h}}_{k}[n]$ is written as [186]
    $\bar{\mathbf{h}}_{k}[n]=\mathrm{e}^{j\theta^{\prime}_{k}}\bar{\mathbf{h}}_{k}[n-1],$ (27)
    where $\bar{\mathbf{h}}_{k}[n-1]$ is the LoS component at time (n1)(n-1). Moreover, $\theta^{\prime}_{k}\in[-\pi,\pi]$ is a phase shift incurred due to user mobility, and it is assumed to be common to all antenna elements when far-field propagation is considered. Next, the effects of channel aging in the non-LoS component can be modeled as [187]
    $\tilde{\mathbf{h}}_{k}[n]=\chi\tilde{\mathbf{h}}_{k}[n-1]+\sqrt{1-\chi^{2}}\mathbf{e}[n],$ (28)
    where $\chi$ is the temporal correlation coefficient, and $\mathbf{e}[n]\sim\mathcal{CN}(\bm{0},\mathbf{I}_{M})$ is the innovation component at the time instance $n$. Here, the temporal correlation coefficient corresponds to the autocorrelation function of Clarke-Gans/Jakes model. It can be defined as $\chi=\mathrm{J}_{0}(2\pi f_{D}T_{s})$, where $\mathrm{J}_{0}(\cdot)$ is the zeroth-order Bessel function of the first kind, $f_{D}=\nu f_{c}/c$ is the maximum Doppler shift, $T_{s}$ is the channel sampling duration, $f_{c}$ is the carrier frequency, and $c$ is the velocity of light.
  2. II.
    Topology Variations: Two system topologies are considered by changing the number of users as $K=4$ and $K=8$. First, the ISAC waveform, which is optimized through our learning model for four users, is used for data transmission when there actually exist eight users. Second, the optimal waveform for eight users is also obtained by retraining the model and used it for data transmission towards those eight users.
The performance will be degraded when the learning-based ISAC waveforms are deployed in real-time system operations due to their generalization limitations in dynamic propagation environments. This is because the inferred ISAC waveforms from the ML models that are trained for the aged channels/topologies will no longer be optimal for dynamic systems with user mobility and volatility in concurrent connections.
Figure 8: Performance degradations of the learning-based ISAC waveforms; channel aging and topology changes. System parameters are set to $M=16$, $K=\{4,8\}$, $\tau_{d}=10$, and $\eta=0.2$.
In Fig. 8, the performance degradations resulting from generalization limitations of the proposed unsupervised learning-based ISAC waveform $(\mathbf{X})$ design due to channel aging and system topology changes are investigated. The achievable sum rate curves are plotted for four cases. In Case-A, the system consists of $K=4$ users, and $\mathbf{X}$ is optimized via the proposed unsupervised learning framework. In Case-B, the impact of channel aging is studied with user mobility with a velocity of $\nu=2$ m/s. To this end, the ISAC waveform is optimized for the aged channels $(\mathbf{h}_{k}[n-1])$ and used it when the true channels are given by $\mathbf{h}_{k}[n]$. The relationship between $\mathbf{h}_{k}[n-1]$ and $\mathbf{h}_{k}[n]$ is captured through our channel model under user mobility in (26)-(28) with user mobility velocity of $\nu=2$ m/s. Fig. 8 clearly unveils that the generalization limitation of the learned ISAC waveform under channel aging degrades the achievable sum rate. For example, when Case-B is compared to Case-A, at $\rho_{d}=6$ dB, a percentage sum rate degradation of 7.62 % can be observed. Moreover, in Case-C, the user mobility velocity is increased to $\nu=10$ m/s, and the ISAC waveform is learned for the aged channels. When Case-C is compared with Case-A at $\rho_{d}=6$ dB, a higher percentage sum rate loss of 10.42 % can be observed, mainly due to increased user mobility velocity of Case-C over Case-B.
Figure 8 also shows the performance degradation arising from system topology changes in terms of the number of users $K\in\{4,8\}$ in Cases D, and E, and F. In Case-D, the ISAC waveform is optimized to $K=8$ users via our learning-based approach. In Case-F, the number of users increases to $K=8$ while the ISAC waveform optimized for four users is still in use at the BS. A comparison between Case-D and Case-F reveals that the sum rate degrades by 55.24 % at $\rho_{d}=6$ dB when the BS adopts the ISAC waveform optimized for four users, while the system topology changes to eight users. In Case-E, the BS uses an ISAC waveform that is optimized to the aged channels under a user mobility with $\nu=2$ m/s with $K=8$. When Case-E is compared with Case-D, the sum rate is degraded by 8.66 % at $\rho_{d}=6$ dB when the BS uses an ISAC waveform optimized to aged channels while the true channels have changed due to mobility scenario with $K=8$. The above sum rate degradations are primarily caused by the mismatched ISAC waveforms that are no longer optimal for suppressing inter-user interference under extreme user mobility and volatility in concurrent connections. Hence, the observations in Fig. 8 reveal that the learning-based ISAC waveform designs are vulnerable to generalization limitations due to channel aging and system topology changes.